Radicals & rational expressions (solve using appropriate method)
sqrt 2x - 4 = x - 14

Answer by fafa(13)

square root 2x-4 = x-14
we have to square both side so the square root will b gone
2x-4= (x-4)^2
2x-4= x^2 - 28x + 196
we can change the side to solve this case easier
x^2 - 28x + 196 = 2x - 4
x^2 - 30 x + 200 = 0
(x-20).(x-10)=0
x-20=0
x=20
or
x-10=0
x=10
so x=20 or x=10